factor by grouping:
8x^2-34x+21
: factor by grouping:
8x^2-34x+21

Answer by checkley77(3397)

8x^2-34x+21
(4x-3)(2x-7)
Question 152455: factor by grouping:
8x^2-34x+21
: factor by grouping:
8x^2-34x+21

Answer by jim_thompson5910(9166)


Looking at the expression 8x^2-34x+21, we can see that the first coefficient is 8, the second coefficient is -34, and the last term is 21.


Now multiply the first coefficient 8 by the last term 21 to get (8)(21)=168.


Now the question is: what two whole numbers multiply to 168 (the previous product) and add to the second coefficient -34?


To find these two numbers, we need to list all of the factors of 168 (the previous product).


Factors of 168:
1,2,3,4,6,7,8,12,14,21,24,28,42,56,84,168
-1,-2,-3,-4,-6,-7,-8,-12,-14,-21,-24,-28,-42,-56,-84,-168


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to 168.
1*168
2*84
3*56
4*42
6*28
7*24
8*21
12*14
(-1)*(-168)
(-2)*(-84)
(-3)*(-56)
(-4)*(-42)
(-6)*(-28)
(-7)*(-24)
(-8)*(-21)
(-12)*(-14)

Now let's add up each pair of factors to see if one pair adds to the middle coefficient -34:


First NumberSecond NumberSum
11681+168=169
2842+84=86
3563+56=59
4424+42=46
6286+28=34
7247+24=31
8218+21=29
121412+14=26
-1-168-1+(-168)=-169
-2-84-2+(-84)=-86
-3-56-3+(-56)=-59
-4-42-4+(-42)=-46
-6-28-6+(-28)=-34
-7-24-7+(-24)=-31
-8-21-8+(-21)=-29
-12-14-12+(-14)=-26



From the table, we can see that the two numbers -6 and -28 add to -34 (the middle coefficient).


So the two numbers -6 and -28 both multiply to 168 and add to -34


Now replace the middle term -34x with -6x-28x. Remember, -6 and -28 add to -34. So this shows us that -6x-28x=-34x.


8x^2+highlight(-6x-28x)+21 Replace the second term -34x with -6x-28x.


(8x^2-6x)+(-28x+21) Group the terms into two pairs.


2x(4x-3)+(-28x+21) Factor out the GCF 2x from the first group.


2x(4x-3)-7(4x-3) Factor out 7 from the second group. The goal of this step is to make the terms in the second parenthesis equal to the terms in the first parenthesis.


(2x-7)(4x-3) Combine like terms. Or factor out the common term 4x-3

---------------------------------------------


Answer:


So 8x^2-34x+21 factors to (2x-7)(4x-3).


Note: you can check the answer by FOILing (2x-7)(4x-3) to get 8x^2-34x+21 or by graphing the original expression and the answer (the two graphs should be identical).